# TDDBench: A Benchmark for Training data detection

### v1

@@ Abstract
Training Data Detection (TDD) is a task aimed at determining whether a specific data instance is used to train a machine learning model. In the computer security literature, TDD is also referred to as Membership Inference Attack (MIA). Given its potential to assess the risks of training data breaches, ensure copyright authentication, and verify model unlearning, TDD has garnered significant attention in recent years, leading to the development of numerous methods. Despite these advancements, there is no comprehensive benchmark to thoroughly evaluate the effectiveness of TDD methods. In this work, we introduce TDDBench, which consists of 13 datasets spanning three data modalities: image, tabular, and text. We benchmark 21 different TDD methods across four detection paradigms and evaluate their performance from five perspectives: average detection performance, best detection performance, memory consumption, and computational efficiency in both time and memory. With TDDBench, researchers can identify bottlenecks and areas for improvement in TDD algorithms, while practitioners can make informed trade-offs between effectiveness and efficiency when selecting TDD algorithms for specific use cases. Our large-scale benchmarking also reveals the generally unsatisfactory performance of TDD algorithms across different datasets. To enhance accessibility and reproducibility, we open-source TDDBench for the research community.
@@ 1 Introduction
To address these limitations, we introduce TDDBench, a comprehensive framework for benchmarking TDD algorithms. Figure 1 provides an overview of TDDBench. The benchmark includes 13 datasets across three data modalities (tabular, text, and image) and evaluates 21 state-of-the-art TDD algorithms on 41 different target models, including the large language model Pythia-12B. We also categorize the 21 TDD algorithms into four types based on their algorithmic characteristics, including metric-based, learning-based, model-based, and query-based. Using this new benchmark, we conduct large-scale experiments to thoroughly assess TDD algorithms. Specifically, we aim to investigate: 1). The performance of TDD algorithms across various datasets and data modalities. 2). The impact of the target model on TDD algorithms. 3). The limitations and areas for improvement in TDD algorithms. 4). The performance of TDD algorithms from multiple perspectives, including detection performance, practicality, and efficiency in terms of time and memory usage.
The large-scale experimental results reveal several key findings. First, there is a significant performance gap between different types of TDD algorithms, with model-based TDD methods generally outperforming the others. However, this outperformance comes at a cost, as model-based methods require building multiple reference models, leading to high computational expenses. Second, memorization of training data plays a crucial role in the performance of TDD algorithms, with larger target models—often prone to memorization—exhibiting higher TDD success rates. Third, the performance of TDD algorithms is highly dependent on knowledge of the underlying target model architecture. Overall, our experiments show that there is no single best method across all scenarios, and notably, many TDD algorithms perform poorly on data modalities beyond images, indicating the need for further improvement in non-image domains.
@@
A novel and comprehensive TDD benchmark: We introduce TDDBench, a benchmark consisting of datasets across three modalities—image, table, and text. We have open-sourced TDDBench for the research community at https://anonymous.4open.science/r/TDDBench-8078.
New insights in TDD performance: By benchmarking 21 state-of-the-art TDD algorithms, we provide insights into recent advancements in TDD, including strategies for reducing reliance on model-specific knowledge and maximizing the benefits of greater computational resources.
@@ 3.2 Overall detection performance across different datasets and models
Overall performance is not satisfactory. In most experimental settings, the AUC scores range between 0.5 and 0.6. From an AUC perspective, this is clearly unsatisfactory, as it indicates a high rate of false negatives and false positives. In other words, data points used by the target model are frequently misclassified as not being used, and vice versa. This is concerning and highlights the urgent need for advancing the performance of TDD methods. For model-based and query-based methods, whose results may be influenced by the number of reference models and queries, we perform a robust analysis by varying the number of reference models and queries. The results remain largely consistent across different configurations, as shown in Appendix A.4.
@@ 3.3.1 Data Memorization and Overfitting
A common view is that the effectiveness of TDD is closely tied to the level of training-data memorization or overfitting exhibited by the target model during training (Yeom et al., 2018; Long et al., 2018). The disparity between the target model’s accuracy on the training set and the test set, known as the train-test accuracy gap, serves as an indicator of data memorization in prior literature (Carlini et al., 2022a). In our experiments, we document the train-test gaps of 12 distinct target models in Table 4, along with the corresponding performance of all detection methods. Specifically, the training of target models is repeated five times with different random training samples. Figure 2 illustrates the performance of various detection algorithms across different train-test gaps, with error bars representing 95% confidence intervals obtained from five independent trials. It is evident that the performance of all TDD methods is positively correlated with the train-test accuracy gap of the target model. When the train-test gap is maximal, indicating the target model may overfit to the training data, identifying whether a data point was used by the target model becomes relatively easier. In contrast, when the target model has a minimal train-test gap, the TDD algorithms often fall short in detection performance. Therefore, it is crucial for future TDD research to examine the generalizability of target models and report TDD performance when the train-test gap is small.
@@ 3.4 Performance when Knowledge about the Target Model is Limited
The results, presented in Table 6, show a noticeable decline in detection performance when the data owner has limited knowledge about the target model’s architecture. For example, Learn-original exhibits a 5.3% performance decline when using ResNet18 as the shadow model. This performance degradation can be attributed to discrepancies between the shadow and target models, which result in biased input features for training the auxiliary classifier. Moreover, the overall AUC performance remains unsatisfactory, highlighting the ineffectiveness of TDD algorithms in most datasets when knowledge about the target model is limited.
@@ 4 Related Work
Training data detection (TDD) is commonly employed to assess privacy risks in machine learning models (Murakonda & Shokri, 2020). It has been applied across various domains, including image classification (Hui et al., 2021; Rezaei & Liu, 2021), text generation (Shejwalkar et al., 2021; Oh et al., 2022), graph neural networks (Wu et al., 2021; Olatunji et al., 2021), and recommendation systems (Zhang et al., 2021; Wang et al., 2022). TDD has a wide range of applications such as dataset copyright protection (Maini et al., 2021) and for verifying machine unlearning (Chen et al., 2021).
Shokri et al., 2017 introduce the first TDD algorithm, utilizing shadow models to help identifying differences in the model’s predictions for training data versus other data. Yeom et al., 2018 demonstrate that satisfactory results could be achieved by utilizing only the loss of the target model on the sample. Building on this foundation, subsequent studies employ the entropy (Song & Mittal, 2021) and maximum confidence (Song et al., 2019) of the target model’s predictions as additional detection criteria. Carlini et al., 2022a criticize methods based solely on the target model’s output, arguing that they overlook the inherent characteristics of the data, which can lead to biased estimates regarding whether a sample belongs to the training dataset. They propose training multiple reference models to better understand how the sample’s characteristics influence metrics like loss. Additionally, increasing the number of queries (Rahimian et al., 2020) to the target model is a viable improvement for TDD algorithms, with newly added query samples typically being neighbors of the data under detection, enriching the detection criterion. In short, there is a rapidly growing body of literature on TDD methods, and we provide a brief summarization in Section 2.3.
@@
5 Conclusion
In this article, we introduce TDDBench, a novel and comprehensive training data detection benchmark. Unlike existing benchmarks, TDDBench extends evaluations across multiple data modalities, including image, tabular, and text. It also implements large-scale target models and benchmarks 21 state-of-the-art TDD algorithms. Our comprehensive evaluation sheds critical light on the development of TDD algorithms and helps both researchers and practitioners reconsider the trade-offs involved in using TDD algorithms. For example, our evaluation shows that model-based TDD algorithms outperform others but at the cost of higher time and memory complexity. Additionally, all existing TDD algorithms experience performance degradation when the target model avoids overfitting. Based on our findings with TDDBench, we believe future work on TDD algorithms should focus on, but not be limited to: (1) designing TDD algorithms robust to target models less prone to overfitting, (2) creating TDD algorithms that require minimal knowledge of the target model architecture, (3) achieving a better balance between performance and practical considerations such as computational complexity, and (4) developing algorithms tailored to specific application contexts or data modalities, such as training data detection for large language models and recommendation systems.
@@ Appendix A Appendix
A.1 Details of the algorithms included in TDDBench
A.1.1 Metric-based dectection

### v2

@@ Abstract
Training Data Detection (TDD) is a task aimed at determining whether a specific data instance is used to train a machine learning model. In the computer security literature, TDD is also referred to as Membership Inference Attack (MIA). Given its potential to assess the risks of training data breaches, ensure copyright authentication, and verify model unlearning, TDD has garnered significant attention in recent years, leading to the development of numerous methods. Despite these advancements, there is no comprehensive benchmark to thoroughly evaluate the effectiveness of TDD methods. In this work, we introduce TDDBench, which consists of 13 datasets spanning three data modalities: image, tabular, and text. We benchmark 21 different TDD methods across four detection paradigms and evaluate their performance from five perspectives: average detection performance, best detection performance, memory consumption, and computational efficiency in both time and memory. With TDDBench, researchers can identify bottlenecks and areas for improvement in TDD algorithms, while practitioners can make informed trade-offs between effectiveness and efficiency when selecting TDD algorithms for specific use cases. Our extensive experiments also reveal the generally unsatisfactory performance of TDD algorithms across different datasets. To enhance accessibility and reproducibility, we open-source TDDBench for the research community at https://github.com/zzh9568/TDDBench.
@@ 1 Introduction
To address these limitations, we introduce TDDBench, a comprehensive framework for benchmarking TDD algorithms. Figure 1 provides an overview of TDDBench. The benchmark includes 13 datasets across three data modalities (tabular, text, and image) and evaluates 21 state-of-the-art TDD algorithms on 41 different target models, including the large language model Pythia-12B. We also categorize the 21 TDD algorithms into four types based on their algorithmic characteristics, including metric-based, learning-based, model-based, and query-based. Using this new benchmark, we conduct extensive experiments to thoroughly assess TDD algorithms. Specifically, we aim to investigate: 1). The performance of TDD algorithms across various datasets and data modalities. 2). The impact of the target model on TDD algorithms. 3). The limitations and areas for improvement in TDD algorithms. 4). The performance of TDD algorithms from multiple perspectives, including detection performance, practicality, and efficiency in terms of time and memory usage.
The experimental results reveal several key findings. First, there is a significant performance gap between different types of TDD algorithms, with model-based TDD methods generally outperforming the others. However, this outperformance comes at a cost, as model-based methods require building multiple reference models, leading to high computational expenses. Second, memorization of training data plays a crucial role in the performance of TDD algorithms, with larger target models—often prone to memorization—exhibiting higher TDD success rates. Third, the performance of TDD algorithms is highly dependent on knowledge of the underlying target model architecture. Overall, our experiments show that there is no single best method across all scenarios, and notably, many TDD algorithms perform poorly on data modalities beyond images, indicating the need for further improvement in non-image domains.
@@
A novel and comprehensive TDD benchmark: We introduce TDDBench, a comprehensive, open-source benchmark for training data detection (TDD), encompassing three data modalities and 21 state-of-the-art TDD algorithms. Our experiments underscore the limitations of current TDD algorithms, which face challenges in handling diverse data modalities and various model architectures.
New insights in TDD performance: By benchmarking 21 state-of-the-art TDD algorithms across various detection scenarios, we identify promising directions for future research in TDD, many of which have not been thoroughly explored in prior benchmarks.
@@ 3.2 Overall detection performance across different datasets and models
Overall performance is not satisfactory. In most experimental settings, the AUC scores range between 0.5 and 0.6. From an AUC perspective, this is clearly unsatisfactory, as it indicates a high rate of false negatives and false positives. In other words, data points used by the target model are frequently misclassified as not being used, and vice versa. This is concerning and highlights the urgent need for advancing the performance of TDD methods.
@@ 3.3.1 Data Memorization and Overfitting
A common view is that the effectiveness of TDD is closely tied to the level of training-data memorization or overfitting exhibited by the target model during training (Yeom et al., 2018; Long et al., 2018). The disparity between the target model’s accuracy on the training set and the test set, known as the train-test accuracy gap, serves as an indicator of data memorization in prior literature (Carlini et al., 2022a). In our experiments, we document the train-test gaps of 12 distinct target models in Table 4, along with the corresponding performance of all detection methods. Specifically, the training of target models is repeated five times with different random training samples. Figure 2 illustrates the performance of various detection algorithms across different train-test gaps, with error bars representing 95% confidence intervals obtained from five independent trials. It is evident that the performance of all TDD methods is positively correlated with the train-test accuracy gap of the target model.
Takeaway. It is crucial for future advanced TDD algorithms to evaluate the generalizability of target models and report TDD performance when the train-test gap is small.
@@ 3.3.2 Target model size
Takeaway. TDD algorithms generally demonstrate improved performance as the model size increases, highlighting their potential in the era of large models.
@@ 3.4 Performance when Knowledge about the Target Model is Limited
The results, presented in Table 6, show a noticeable decline in detection performance when the data owner has limited knowledge about the target model’s architecture. For example, Learn-original exhibits a 5.3% performance decline when using ResNet18 as the shadow model. This performance degradation can be attributed to discrepancies between the shadow and target models, which result in biased input features for training the auxiliary classifier.
Takeaway. The overall performance of TDD algorithms, without knowledge of the target model’s training algorithm, remains unsatisfactory. This underscores the ineffectiveness of TDD algorithms on most datasets when information about the target model is limited.
@@ 3.5 Performance tradeoff of TDD algorithms
Takeaway. None of the TDD algorithms are satisfactory, as performance improvements often necessitate increased consumption of computing resources.
@@ 4 Related Work
Training data detection (TDD) is commonly employed to assess privacy risks in machine learning models (Murakonda & Shokri, 2020). It has been applied across various domains, including image classification (Hui et al., 2021), text generation (Shejwalkar et al., 2021), graph neural networks (Wu et al., 2021), and recommendation systems (Zhang et al., 2021). TDD has a wide range of applications such as dataset copyright protection (Maini et al., 2021) and for verifying machine unlearning (Chen et al., 2021). Shokri et al., 2017 introduce the first TDD algorithm, utilizing shadow models to help identifying differences in the model’s predictions for training data versus other data. Yeom et al., 2018 demonstrate that satisfactory results could be achieved by utilizing only the loss of the target model on the sample. Carlini et al., 2022a criticize methods based solely on the target model’s output, arguing that they overlook the inherent characteristics of the data, which can lead to biased estimates regarding whether a sample belongs to the training dataset. They propose training multiple reference models to better understand how the sample’s characteristics influence metrics like loss. There is a rapidly growing body of literature on TDD methods, and we provide a brief summarization in Section 2.3.
@@
5 Conclusions
In this article, we introduce TDDBench, a novel and comprehensive training data detection benchmark. Unlike existing benchmarks, TDDBench extends evaluations across multiple data modalities, including image, tabular, and text. It also includes large language models and benchmarks 21 state-of-the-art TDD algorithms. Our comprehensive evaluation sheds critical light on the development of TDD algorithms and helps both researchers and practitioners reconsider the trade-offs involved in using TDD algorithms. Based on our findings with TDDBench, we believe future work on TDD algorithms should focus on, but not be limited to: (1) designing TDD algorithms robust to target models less prone to overfitting, (2) creating TDD algorithms that require minimal knowledge of the target model architecture, (3) achieving a better balance between performance and practical considerations such as computational complexity, and (4) developing algorithms tailored to specific application contexts or data modalities, such as training data detection for large language models and recommendation systems.
6 ACKNOWLEDGEMENTS
The work was supported by grants from the National Natural Science Foundation of China (No. U24A20253).
@@ Appendix A Appendix
A.1 Training data detection on vision transformer models
In this section, we showcase the performance of the TDD algorithm on vision transformer models. Specifically, we trained ViT (Dosovitskiy, 2020) and Swin (Liu et al., 2021) models on the CIFAR-10 dataset and evaluated the TDD algorithm’s effectiveness on these models. As shown in Table 8 and Table 9, the TDD algorithm remains effective on vision transformer models, and model-based algorithms continue to have clear advantages over other types of methods.
Table 8: TDD performance across different metrics on ViT (Dosovitskiy, 2020) trained on CIFAR-10 dataset. MA(membership advantage) (Jayaraman et al., 2021) equals the difference between the true positive rate and the false positive rate. For all metrics except for FPR and FNR, higher values indicate better performance of the corresponding TDD algorithm.
Metric-loss 0.555 0.811 0.659 0.583 0.190 0.644 0.167 0.010 0.119 0.599
Metric-conf 0.555 0.811 0.659 0.583 0.190 0.644 0.167 0.010 0.119 0.599
Metric-corr 0.555 0.796 0.654 0.581 0.204 0.633 0.162 0.000 0.000 0.581
Metric-ent 0.536 0.499 0.517 0.536 0.501 0.428 0.071 0.011 0.114 0.543
Metric-ment 0.555 0.810 0.659 0.583 0.190 0.644 0.167 0.011 0.121 0.599
Learn-original 0.522 0.675 0.589 0.532 0.325 0.612 0.063 0.013 0.122 0.537
Learn-top3 0.539 0.477 0.506 0.536 0.523 0.405 0.072 0.010 0.116 0.541
Learn-sorted 0.539 0.474 0.504 0.536 0.526 0.402 0.072 0.010 0.116 0.541
Learn-label 0.552 0.787 0.649 0.577 0.213 0.634 0.153 0.015 0.141 0.597
Learn-merge 0.556 0.799 0.655 0.583 0.201 0.634 0.165 0.016 0.138 0.604
Model-loss 0.596 0.718 0.651 0.617 0.283 0.483 0.235 0.056 0.244 0.672
Model-calibration 0.578 0.764 0.658 0.606 0.236 0.553 0.212 0.046 0.222 0.653
Model-lira 0.562 0.735 0.637 0.584 0.265 0.567 0.168 0.052 0.219 0.631
Model-fpr 0.611 0.594 0.603 0.610 0.406 0.374 0.220 0.036 0.231 0.651
Model-robust 0.603 0.666 0.633 0.615 0.334 0.436 0.231 0.056 0.255 0.671
Query-augment 0.557 0.771 0.646 0.581 0.229 0.609 0.162 0.003 0.067 0.594
Query-transfer 0.525 0.738 0.614 0.538 0.262 0.662 0.077 0.009 0.102 0.538
Query-adv 0.564 0.793 0.659 0.593 0.207 0.608 0.186 0.008 0.119 0.590
Query-neighbor 0.504 0.393 0.442 0.505 0.607 0.383 0.010 0.001 0.061 0.506
Query-qrm 0.555 0.801 0.656 0.583 0.199 0.636 0.165 0.000 0.000 0.597
Query-ref 0.676 0.558 0.611 0.649 0.442 0.260 0.298 0.089 0.308 0.718
Table 9: TDD performance across different metrics on Swin (Liu et al., 2021) trained on CIFAR-10 dataset. MA(membership advantage) (Jayaraman et al., 2021) equals the difference between the true positive rate and the false positive rate. For all metrics except for FPR and FNR, higher values indicate better performance of the corresponding TDD algorithm.
Metric-loss 0.571 0.855 0.685 0.609 0.146 0.636 0.218 0.009 0.116 0.621
Metric-conf 0.571 0.855 0.685 0.609 0.146 0.636 0.218 0.009 0.116 0.621
Metric-corr 0.560 0.915 0.695 0.601 0.085 0.713 0.202 0.000 0.000 0.601
Metric-ent 0.547 0.690 0.611 0.562 0.310 0.566 0.124 0.009 0.115 0.573
Metric-ment 0.571 0.856 0.685 0.609 0.144 0.638 0.218 0.009 0.117 0.621
Learn-original 0.548 0.685 0.609 0.563 0.315 0.559 0.126 0.010 0.132 0.577
Learn-top3 0.553 0.625 0.587 0.562 0.375 0.501 0.124 0.009 0.116 0.573
Learn-sorted 0.553 0.624 0.586 0.562 0.376 0.500 0.124 0.009 0.116 0.573
Learn-label 0.571 0.857 0.686 0.610 0.143 0.638 0.219 0.024 0.169 0.638
Learn-merge 0.574 0.838 0.681 0.610 0.162 0.618 0.220 0.020 0.172 0.643
Model-loss 0.623 0.693 0.656 0.639 0.307 0.416 0.278 0.076 0.275 0.704
Model-calibration 0.581 0.810 0.676 0.615 0.190 0.580 0.230 0.061 0.224 0.668
Model-lira 0.590 0.772 0.669 0.621 0.228 0.531 0.241 0.083 0.278 0.686
Model-fpr 0.615 0.636 0.625 0.621 0.364 0.394 0.242 0.058 0.275 0.670
Model-robust 0.606 0.760 0.674 0.635 0.240 0.489 0.271 0.083 0.300 0.705
Query-augment 0.567 0.847 0.679 0.603 0.153 0.640 0.206 0.009 0.019 0.618
Query-transfer 0.542 0.850 0.662 0.570 0.150 0.711 0.139 0.009 0.105 0.560
Query-adv 0.577 0.932 0.713 0.627 0.068 0.677 0.254 0.008 0.155 0.645
Query-neighbor 0.505 0.422 0.460 0.506 0.578 0.410 0.012 0.002 0.079 0.507
Query-qrm 0.567 0.866 0.686 0.606 0.134 0.655 0.211 0.000 0.000 0.623
Query-ref 0.639 0.842 0.726 0.689 0.158 0.464 0.378 0.128 0.383 0.759
A.2 Training data detection algorithms in large language models
In this section, we offer a brief introduction to the TDD algorithms for large language models; for more detailed information, please refer to the related works.
MIN-K% (Shi et al., 2024) utilizes the k% of tokens with the lowest likelihoods as the detection basis, rather than average loss.
PAC, short for Polarized Augment Calibration (Ye et al., 2024), introduces a new detection metric called polarized distance through data augmentation. This metric helps determine whether data has been trained by large language models.
Additionally, we present the performance of the TDD algorithm on various sizes of Llama models (Touvron et al., 2023). The experimental results in Table 10 indicate that the performance of the TDD algorithm improves as the model size increases, which aligns with the results observed for the detection results on Pythia (corresponds to Figure 4 in Section 3.3.2) .
Table 10: The performance of the TDD algorithm across different sizes of Llama models.
Algorithm Target Model
Llama-7b Llama-13b Llama-30b Llama-65b
Neighbor 0.555 0.552 0.566 0.586
LOSS 0.666 0.678 0.704 0.707
PAC 0.679 0.689 0.704 0.714
Zlib 0.683 0.697 0.718 0.721
MIN-K% 0.697 0.715 0.737 0.737
Reference 0.802 0.809 0.833 0.831
A.3 Training data detection in semi-supervised learning scenarios
Research on TDD algorithms has primarily concentrated on two types of training methods. The first is supervised learning, which forms the basis for most TDD algorithms, covering various fields such as image (Carlini et al., 2022a), table (Shokri et al., 2017), and text (Amit et al., 2024). This is also the setting for our main experiments. The second type is self-supervised learning, which typically focuses on detecting whether the pre-trained corpus of large language models can be identified. This type of algorithm is also known as pretraining data detection (Shi et al., 2024), and our experiments on Llama and Pythia evaluated the TDD algorithm’s performance in this setting.
Table 11: TDD performance on WRN28-2 trainied with semi-supervised method FixMatch (Sohn et al., 2020).
Dataset CIFAR-10 CIFAR-100
Algorithm Accuracy Precision Recall F1-score Accuracy Precision Recall F1-score
Metric-loss 0.624 0.560 0.805 0.661 0.714 0.598 0.721 0.653
Metric-conf 0.624 0.560 0.805 0.661 0.714 0.598 0.721 0.653
Metric-corr 0.533 0.487 0.910 0.635 0.683 0.531 0.786 0.634
Metric-ent 0.643 0.567 0.870 0.687 0.736 0.587 0.820 0.684
Metric-ment 0.624 0.561 0.804 0.661 0.714 0.602 0.712 0.652
Learn-original 0.641 0.564 0.889 0.690 0.737 0.585 0.832 0.687
Learn-top3 0.643 0.569 0.862 0.685 0.733 0.591 0.801 0.680
Learn-sorted 0.643 0.568 0.865 0.686 0.734 0.591 0.804 0.681
Learn-label 0.617 0.550 0.835 0.664 0.722 0.584 0.778 0.668
Learn-merge 0.621 0.560 0.789 0.655 0.717 0.607 0.712 0.655
Model-loss 0.620 0.561 0.775 0.651 0.721 0.606 0.726 0.661
Model-calibration 0.604 0.546 0.774 0.641 0.690 0.555 0.737 0.633
Model-lira 0.614 0.555 0.778 0.648 0.720 0.628 0.685 0.655
Model-fpr 0.597 0.555 0.664 0.605 0.694 0.613 0.625 0.619
Model-robust 0.589 0.526 0.868 0.655 0.726 0.605 0.745 0.668
Query-augment 0.575 0.515 0.888 0.652 0.694 0.576 0.697 0.631
Query-transfer 0.531 0.496 0.589 0.539 0.595 0.453 0.699 0.550
Query-adv 0.607 0.561 0.753 0.643 0.749 0.649 0.769 0.704
Query-neighbor 0.505 0.473 0.582 0.522 0.515 0.393 0.470 0.428
Query-qrm 0.611 0.562 0.710 0.628 0.697 0.595 0.671 0.631
Query-ref 0.642 0.540 0.735 0.623 0.738 0.548 0.711 0.619
In this part, we conduct experiments on the CIFAR-10 and CIFAR-100 datasets to assess the effectiveness of TDD algorithms in semi-supervised learning scenarios, as opposed to the more commonly studied supervised learning setting. To this end, we evaluate the performance of TDD on WRN28-2 trained using FixMatch (Sohn et al., 2020), a state-of-the-art semi-supervised learning method. This investigation aims to provide insights into the versatility and robustness of TDD algorithms in diverse learning settings.
The experimental results lead to the following conclusions: The TDD detection method remains effective in semi-supervised training, but its performance declines compared to supervised learning. Specifically, the model-based method, which shows clear advantages in supervised learning, performs moderately in the semi-supervised setting. This may be because the model-based approach relies on training a reference model, and its performance is significantly affected when the reference model is unaware of the semi-supervised learning method used by the target model.
Future direction. Based on the experimental findings, we believe that investigating TDD algorithms tailored to specific training methods represents a promising and impactful research direction.
A.4 Defense strategies against training data detection
In the field of computer security, training data detection is known as a Membership Inference Attack, which aims to extract private information about the training data from target models. To counteract this detection, various measures (Baek & Shim, 2022; Ying et al., 2020) have been proposed. Since the effectiveness of training data detection is often linked to the degree of overfitting in the target model, many defense methods focus on reducing overfitting. These methods include dropout strategies (Salem et al., 2019), label smoothing (Kaya & Dumitras, 2021), early stopping (Song & Mittal, 2021), and data augmentation (Kaya & Dumitras, 2021).
Beyond reducing model overfitting, another key defense strategy involves modifying the output vector of the target model to lower the risk of training data leakage. For instance, Jia et al., 2019 suggests adding carefully designed noise to the model’s output vector, which does not affect the target model’s performance but can mislead detection algorithms. Shokri et al., 2017 recommends constraining the target model to output only prediction labels without confidence scores, rendering many TDD algorithms ineffective.
In addition to these common methods applicable across different data types, Hayes et al., 2017 employs differential privacy to prevent external parties from determining whether specific data was used in a generative model. Their results indicate that differential privacy can balance model usability with defense effectiveness. Shejwalkar et al., 2021 proposed a defense strategy based on knowledge distillation, demonstrating that the distilled model can better resist training data detection.
To better assess the robustness of TDD algorithms, we examine their performance when the target model is combined with various defense strategies. Specifically, we selected four general defense strategies: using dropout and label smoothing on the target model to mitigate overfitting, and altering the target model’s output vector to noise vectors and hard label. Our experimental results across three datasets indicate that these defense strategies, particularly the addition of noise, can effectively diminish the performance of TDD algorithms. TDD algorithms with strong performance, such as those that are learning-based and model-based, heavily depend on the authenticity of the model’s output vectors. Introducing small amounts of noise to the model output can significantly compromise the effectiveness of these TDD algorithms.
Future direction. Based on these findings, we suggest that to counter potential defense mechanism of the target model, a promising direction is to develop adaptive TDD approaches, which involve designing more effective TDD algorithms tailored to specific defense strategies.
Table 12: TDD performance under different defense strategies.
Dataset CIFAR10(Image) Purchase(Tabular) Rotten-tomatoes(Text)
Defense None Dropout Smooth Noise Label-only None Dropout Smooth Noise Label-only None Dropout Smooth Noise Label-only
Metric-loss 0.635 0.557 0.589 0.558 N/A 0.620 0.615 0.657 0.623 N/A 0.582 0.577 0.590 0.587 N/A
Metric-conf 0.635 0.557 0.589 0.558 N/A 0.620 0.615 0.657 0.623 N/A 0.582 0.577 0.590 0.587 N/A
Metric-corr 0.552 0.547 0.557 0.549 0.552 0.551 0.558 0.556 0.551 0.551 0.557 0.546 0.555 0.554 0.557
Metric-ent 0.628 0.543 0.568 0.543 N/A 0.616 0.606 0.655 0.620 N/A 0.561 0.560 0.573 0.570 N/A
Metric-ment 0.635 0.557 0.589 0.558 N/A 0.620 0.615 0.656 0.623 N/A 0.582 0.577 0.590 0.587 N/A
Learn-original 0.631 0.540 0.493 0.539 N/A 0.652 0.617 0.605 0.659 N/A 0.558 0.558 0.574 0.572 N/A
Learn-top3 0.628 0.543 0.593 0.543 N/A 0.677 0.616 0.652 0.680 N/A 0.561 0.560 0.573 0.570 N/A
Learn-sorted 0.628 0.543 0.576 0.543 N/A 0.667 0.615 0.651 0.678 N/A 0.561 0.560 0.573 0.570 N/A
Learn-label 0.634 0.557 0.565 0.554 N/A 0.656 0.629 0.635 0.662 N/A 0.584 0.578 0.591 0.588 N/A
Learn-merge 0.656 0.559 0.609 0.559 N/A 0.684 0.630 0.664 0.690 N/A 0.584 0.577 0.593 0.592 N/A
Model-loss 0.664 0.606 0.614 0.607 N/A 0.725 0.693 0.732 0.726 N/A 0.773 0.696 0.677 0.731 N/A
Model-calibration 0.639 0.596 0.582 0.592 N/A 0.684 0.671 0.717 0.685 N/A 0.695 0.647 0.624 0.665 N/A
Model-lira 0.690 0.564 0.561 0.568 N/A 0.755 0.698 0.703 0.756 N/A 0.753 0.683 0.678 0.721 N/A
Model-fpr 0.647 0.571 0.518 0.568 N/A 0.697 0.651 0.547 0.703 N/A 0.679 0.653 0.537 0.675 N/A
Model-robust 0.635 0.611 0.642 0.613 N/A 0.711 0.692 0.761 0.712 N/A 0.766 0.692 0.708 0.730 N/A
Query-augment 0.511 0.551 0.586 0.547 0.511 0.533 0.602 0.619 0.612 0.533 0.533 0.559 0.571 0.571 0.533
Query-transfer 0.573 0.534 0.522 0.524 0.573 0.612 0.533 0.534 0.529 0.612 0.570 0.531 0.521 0.526 0.570
Query-adv 0.522 0.548 0.591 0.539 0.522 0.529 0.607 0.629 0.623 0.529 0.530 0.559 0.571 0.570 0.530
Query-neighbor 0.615 0.494 0.481 0.515 N/A 0.620 0.532 0.543 0.534 N/A 0.571 0.523 0.537 0.529 N/A
Query-qrm 0.532 0.575 0.633 0.570 N/A 0.523 0.617 0.658 0.650 N/A 0.524 0.574 0.590 0.589 N/A
Query-ref 0.735 0.604 0.750 0.497 N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
A.5 Training data detection with different sizes of data points
To examine the effect of the size of evaluated data points on the TDD algorithm, we varied the size of the target dataset and assessed the algorithm’s performance on target models trained with these different dataset sizes. As shown in Figure 1, the model-based method consistently delivers the best detection performance across various sizes, which aligns with earlier findings. Furthermore, there is no strong correlation between data size and the performance of the TDD algorithm.
(a) Image:CIFAR10
(b) Tabular:Purchase
(c) Text:Rotten-tomatoe
Figure 4: TDD algorithm performance (AUROC) versus data size, measured by the number of data points in the target dataset.
A.6 Details of the algorithms included in TDDBench
A.6.1 Metric-based dectection